OpenStudy (anonymous):

solve the equation for x: sqrt{x-5}=x-7

jhonyy9 (jhonyy9):

x*2 -10x +25 = x-7 --- if you have thought (x-5)*(2) where 2 is exponent

OpenStudy (angela210793):

sqrt{x-5}=x-7(Raise in the 2nd power both sides and u'll have: x-5=x^2-14x+49 x^2-15x+54=0 (x-9)*(x-6)=0 x=9 or x=6

OpenStudy (angela210793):

actually now tht i'm seeing it the solution is only 9 not 6